OpenStudy (anonymous):

Two students standing apart in a classroom have a gravitational force of attraction between them of 3.20 x 10-8 N. If one student has a mass of 50.0 kg and the other has a mass of 60.0 kg, what is the distance between the students?

OpenStudy (anonymous):

I really need help

OpenStudy (isaiah.feynman):

You will need this law of universal gravitation..\[F = G \frac{ m_{1}m_{2} }{ r^{2} }\]

OpenStudy (isaiah.feynman):

You want to know r so we can tweak the formula a bit \[r = \sqrt{G\frac{ m_{1}m_{2} }{ F }}\]